The common stock of the P.U.T.T. Corporation has been trading in a narrow price range for the past month, and you are convinced it is going to break far out of that range in the next three months. You do not know whether it will go up or down, however. The current price of the stock is $115 per share, the price of a three-month call option with an exercise price of $115 is $8, and a put with the same expiration date and exercise price costs $6.


a. What would be a simple options strategy to exploit your conviction about the stock price’s future movements?

  • Sell both a put option and a call option on the stock.

  • Buy both a put option and a call option on the stock.

  • Buy a put option and sell a call option on the stock.

  • Sell a put option and buy a call option on the stock.

b. How far would the price have to move in either direction for you to make a profit on your initial investment?

Stock price must move up above $________, or move down below $_______for you to make a profit
Answer #1

As the stock can give break out either direction, you should both a call and a put option. (Long Sraddle).

a. Buy both a put option and a call option on the stock.

b. Your investment for buying a call and a put option=8+6=$14

Hence, it should either go above =115+14=$129 or down below 115-14=101 for making profit.

Stock price must move up above $129, or move down below $101 for you to make a profit